find the domain of 2x+3=-3(x-1)
Answer by edjones(8007)

2x+3=-3(x-1)
2x+3=-3x+3
2x+3x+3-3=-3x+3x+3-3
5x=0
5x/5=0/5
x=0
The domain of x=0.
.
Ed